Accessibility Assistance

Skip to Content




Information Technology
George A. Smathers Libraries
University of Florida
P.O. Box 117000
Gainesville, FL 32611-7000

Telephone: (352) 273-2800
Fax: (352) 392-2354

Information Technology Staff Meeting Notes
January 10, 2012

Present: Thomas Bielicke, Cynthia Bowker, Chris Case, Will Chaney, Logan Clapp, Debra Fetzer, Winston Harris, Michael Jay, Amy Polk, Mark Sullivan, Evan Wack
Announcements
  1. Web Developer position extended to January 17th.
  2. Professional Development conversation with Rachel Schipper scheduled for Tuesday, January 17th. Have your spreadsheets done by then.
  3. Darlene - coffee shop lady. Her husband passed over the holiday break. Michael is suggesting we purchase flowers for her.
  4. Chris Nicolich's baby was born last week: Lauren Renae.
  5. Amy's daughter is turning 16.
Ongoing IT Updates

Hardware
  1. App Streaming server patched. Logan spun up three servers for testing. Packaging Acrobat X, Contribute and Captivate are still in testing.
  2. Migration - West laptop cart deployed. The switch did not pick up policy at first, but this has been fixed. MSL laptop cart switch is waiting on CNS to configure. Still plowing through migration reports.
  3. AV - Mike is out sick. Looking for new mics for 1A.
  4. VDI - meeting scheduled for Thursday with AT.
  5. Library West 24/5 - Smart Control issue - computers went into maintenance control the first night. Eventually, the maintenance window was shut off and patron log on was enabled. Hardware unit to be trained on Smart Control.
Software
  1. Kuali OLE - Chris met with Judy before break to discuss the road map for the development cycle of OLE. She requested more information about plans after Release 1. Wanted to know how Single BIB records and UBorrow would fit into the scheme. Chris also met with Enterprise folks for a preliminary action plan on implementation.
  2. Print Journal Inventory Tracker - on hold until Chris Nicolich returns.
  3. ASERL - Tom is working on it.
  4. Circulation switch to "Library Guest" accounts - no movement. AD folks are looking into the issue.
  5. Macros - five open as a result of migration.
  6. ARES-Sakai authentication - trying to get this restarted after break.
  7. Broadcast to Public Computers tool - it is now standalone. Need to implement tracking and handling.
Network/Server Administration
  1. VDI testing - new licenses added.
  2. New web infrastructure - most staff mapped are mapped to the new web folder W: We are taking advantage of access based enumeration for simplicity.
  3. New media server online.
  4. New server requested for the security cameras in East and West.
  5. SharePoint - rollout v2 - planning how to make SharePoint more attractive to staff.
Web & Digital
  1. Web Projects -Web server upgrade - LibWeb2 this week.
  2. Contribute - still making corrections.
  3. ALADN - was off, but now have to recreate an old web page.
  4. On hold: LAC website update, Panama Canal Museum - wall of honor migration (March 1st), ALF/Storage (Cathy), Records Management migration
  5. DLC Projects - FCLA implementation of SobekCM.
  6. New version of open-source SobekCM Web App being released.
  7. DLC-to-Hathi Trust format and workflow tools (on hold)